OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places. X^2+8x+16=1 What are the exact solutions?

jimthompson5910 (jim_thompson5910):

x^2+8x+16=1 (x+4)^2=1 x+4=sqrt(1) x+4=1 or x+4=-1 x=1-4 or x=-1-4 x=-3 or x=-5

OpenStudy (anonymous):

Very well put!